# HOLISMOKES XVI: Lens search in HSC-PDR3 with a neural network committee and post-processing for false-positive removal

### v1

@@ Acknowledgements.
SS has received funding from the European Union’s Horizon 2022 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101105167 — FASTIDIoUS. We acknowledge financial support through grants PRIN-MIUR 2017WSCC32 and 2020SKSTHZ. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (LENSNOVA: grant agreement No 771776). This research is supported in part by the Excellence Cluster ORIGINS which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 – 390783311. This work uses the following software packages: Astropy (Astropy Collaboration et al., 2013; Price-Whelan et al., 2018), matplotlib (Hunter, 2007), NumPy (van der Walt et al., 2011; Harris et al., 2020), Python (Van Rossum & Drake, 2009), Scipy (Virtanen et al., 2020).References
@@
- Acevedo Barroso et al. (2024) Acevedo Barroso, J. A., O’Riordan, C. M., Clément, B., et al. 2024, arXiv e-prints, arXiv:2408.06217
- Andika et al. (2024) Andika, I. T., Schuldt, S., Suyu, S. H., et al. 2024, arXiv e-prints, arXiv:2412.12709
@@
- Enzi et al. (2024) Enzi, W. J. R., Krawczyk, C. M., Ballard, D. J., & Collett, T. E. 2024, arXiv e-prints, arXiv:2411.08565
@@
- Lange et al. (2024) Lange, S. C., Amvrosiadis, A., Nightingale, J. W., et al. 2024, arXiv e-prints, arXiv:2410.12987
@@
- Wong et al. (2020) Wong, K. C., Suyu, S. H., Chen, G. C. F., et al. 2020, MNRAS
@@
Appendix A Color-composite images of re-discovered lens candidates
@@
Appendix B Color-composite images of lens candidates with their predicted mass model
In this section, we show the color-composite image stamps of the lens candidates in analogy to Figs. 6 and 7, as well as those in Sect. A. Contrary to previous figures, we show here the lens center and Einstein radius for each system predicted by the ResNet of Schuldt et al. (2023a). These models were also shown in one panel during the visual inspection to help the grader classify.

### v2

@@ Acknowledgements.
We thank the anonymous referee for comments that helped to improve the clarity of the paper. SS has received funding from the European Union’s Horizon 2022 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101105167 — FASTIDIoUS. We acknowledge financial support through grants PRIN-MIUR 2017WSCC32 and 2020SKSTHZ. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (LENSNOVA: grant agreement No 771776). This research is supported in part by the Excellence Cluster ORIGINS which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 – 390783311. This work uses the following software packages: Astropy (Astropy Collaboration et al., 2013; Price-Whelan et al., 2018), matplotlib (Hunter, 2007), NumPy (van der Walt et al., 2011; Harris et al., 2020), Python (Van Rossum & Drake, 2009), Scipy (Virtanen et al., 2020).References
@@
- Acevedo Barroso et al. (2025) Acevedo Barroso, J. A., O’Riordan, C. M., Clément, B., et al. 2025, A&A, 697, A14
- Andika et al. (2025) Andika, I. T., Schuldt, S., Suyu, S. H., et al. 2025, A&A, 694, A227
@@
- Enzi et al. (2025) Enzi, W. J. R., Krawczyk, C. M., Ballard, D. J., & Collett, T. E. 2025, MNRAS, 540, 247
@@
- Ertl et al. (2025) Ertl, S., Suyu, S. H., Schuldt, S., et al. 2025, arXiv e-prints, arXiv:2503.09718
- Euclid Collaboration et al. (2025a) Euclid Collaboration, Holloway, P., Verma, A., et al. 2025a, arXiv e-prints, arXiv:2503.15328
- Euclid Collaboration et al. (2025b) Euclid Collaboration, Walmsley, M., Holloway, P., et al. 2025b, arXiv e-prints, arXiv:2503.15324
- Frye et al. (2024) Frye, B. L., Pascale, M., Pierel, J., et al. 2024, ApJ, 961, 171
@@
- Lange et al. (2025) Lange, S. C., Amvrosiadis, A., Nightingale, J. W., et al. 2025, MNRAS, 539, 704
@@
- Rodney et al. (2021) Rodney, S. A., Brammer, G. B., Pierel, J. D. R., et al. 2021, Nature Astronomy, 5, 1118
@@
- Wong et al. (2020) Wong, K. C., Suyu, S. H., Chen, G. C. F., et al. 2020, MNRAS, 498, 1420
@@
Appendix A Detailed description of the released catalog of inspected lens candidates
With this publication, we have release the full catalog of inspected lens candidates. This catalog is introduced in Table 2 and electronically available at the HOLISMOKES webpage666www.holismokes.org, the SuGOHI database777https://www-utap.phys.s.u-tokyo.ac.jp/˜oguri/sugohi/, and CDS. The columns are described in Table 3.
| Column | Header | Description |
| (1) | Name | name for grade A and B lens candidates |
| (2) | RA [deg] | Right Ascension of the inspected candidate |
| (3) | Dec [deg] | Declination of the inspected candidate |
| (4) | p | average score predicted by the network committee |
| (5) | G | average grade between 0 and 3 obtained through visual inspection (see Sect. 4) |
| (6) | sigma_G | standard deviation of obtained visual inspection grades |
| (7) | N_graders | number of visual inspectors (see Sect. 4) |
| visual inspection grade from our earlier work S25 | ||
| visual inspection grade from our earlier work Shu et al. (2022) | ||
| visual inspection grade from our earlier work Cañameras et al. (2021) | ||
| (17) | x_err | 1 value for the -center coordinate predicted by the modeling network from S23a |
| (19) | y_err | 1 value for the -center coordinate predicted by the modeling network from S23a |
| (21) | ex_err | 1 value for the -component of the complex ellipticity predicted by the modeling network from S23a |
| (23) | ey_err | 1 value for the -component of the complex ellipticity predicted by the modeling network from S23a |
| (25) | rE_err | 1 value of the Einstein radius predicted by the modeling network from S23a |
| (27) | gma1_err | 1 value of the -component predicted by the modeling network from S23a |
| (29) | gam2_err | 1 value of the -component predicted by the modeling network from S23a |
| (31) | z | photometric redshift value from the catalog compiled by S25 |
| based on DEmP (Hsieh & Yee 2014), Mizuki (Tanaka et al. 2018), and NetZ (Schuldt et al. 2021) | ||
| (36) | Ntot | Sum of objects with photo- within a box of 200″on a side. |
| http://sunguoyou.lamost.org/glc.html.). |
Appendix B Color-composite images of rediscovered lens candidates
@@
Appendix C Color-composite images of lens candidates with their predicted mass model
In this section, we show the color-composite image stamps of the lens candidates in analogy to Figs. 6 and 7, as well as those in Sect. B. Contrary to previous figures, we show here the lens center and Einstein radius for each system predicted by the ResNet of Schuldt et al. (2023a). These models were also shown in one panel during the visual inspection to help the grader classify.
Appendix D Comparison between inspected images from S25 and this work
In this section, we show a representative sample of lens candidates that were graded in this work (see Sect. 4) and in S25. While we showed during grading three different scalings, we show here only one filter combination for simplicity.